# Fast Radio Bursts as a probe of gravity on cosmological scales

### v1

@@ 3.1 MCMC likelihood
3.2 Mock catalogue specification
3.2.1 FRB catalogue
We model the FRB catalogue properties according to a survey that is feasible within the next decade based on the predictions from Petroff et al. (2019). For this analysis, we choose $10^{4}$ FRB samples with associated redshift, whereas a substantial fraction is located above redshift 2. Assuming the FRBs trace the galaxy distribution (Reischke et al., 2021), the source distribution can be written as (assuming a flux limited sample)
$n(z)\propto z^{2}\exp(-\alpha z)\;,$ (32)
where $\alpha$ denotes the survey depth. Henceforth, we use $\alpha=2.5$ which fulfils the demands above. Note that the current sample of the CHIME survey roughly follows this shape with $\alpha=4$. We split the redshift distribution into four equipopulated redshift bins (see also Tanidis & Camera 2019), thus recovering depths information and increasing the signal-to-noise ratio (SNR). We estimate the SNR from the Fisher information matrix (Tegmark et al., 1997) by treating the signal amplitude $A$ defined by $\hat{\boldsymbol{C}}_{\ell}=A\hat{\boldsymbol{S}}_{\ell}+\hat{\boldsymbol{N}}_{\ell}$ as a cosmological parameter, in which case the cumulative SNR can be calculated as
@@ 3.2.1 FRB catalogue
$\Omega_{\mathrm{b}}$ $0.044^{+0.006}_{-0.010}$ $0.049^{+0.010}_{-0.010}$ $0.0482^{+0.003}_{-0.003}$ 3.3 (3.3)
$\Omega_{\mathrm{cdm}}$ $0.240\pm 0.023$ $0.259\pm 0.012$ $0.264\pm 0.003$ 4 (4)
@@ 3.3 Three dimensional power spectrum generation
We calculate the linear matter power spectrum with the Boltzmann solver Horndeski in class (hi_class [^10] http://miguelzuma.github.io/hi_class_public/, Zumalacárregui et al., 2017, Bellini et al., 2020) suite which is an extension to the Cosmic Linear Anisotropy Solving System code (class [^11] https://github.com/lesgourg/class_public, Blas et al., 2011). Subsequently, utilising pyccl [^12] https://github.com/LSSTDESC/CCL (Chisari et al., 2019), we pass the linear matter power spectrum into HMcode (Mead et al., 2015) to calculate the non-linear matter and electron power spectrum $P_{\mathrm{mm}}$, $P_{\mathrm{\mathrm{ee}}}$ and, consequently, the electron bias via
@@
in Planck cosmology. The majority of signal resides around $\ell\sim\mathcal{O}(200)$ (see also Fig. 2) for FRBs. While cosmic shear obtains a significant signal from larger $\ell$, contributions for $k>50\;\mathrm{Mpc}$ will generally be small. For more details on the emulator choices we refer to appendix A.1.
@@ 4 Results
Further peculiar behaviour can be seen for the $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$-Euclid-CS posterior in Fig. 5; displaying a bimodal distribution. While a Gaussian probability distribution cannot be expected in an MCMC, specifically the modified gravity parameter contour shapes show large deviations thereof caused by limiting prior ranges. We excluded significant influence from the MCMC-algorithm as the underlying cause (see App. B). Considering the strong deviation of the predicted $k_{\mathrm{s}}$ to the fiducial one, we conclude that the contours from Fig. 5 to be strongly influenced by the informative prior we place on $\hat{\alpha}_{\mathrm{B/M}}$ and $k_{\mathrm{s}}$.
@@
The motivation behind this choice is of both physical and numerical nature. In principle, $\hat{\alpha}_{\mathrm{B/M}}$ can be negative (Bellini & Sawicki, 2014) and even are expected to in many of the commonly investigated scalar-field theories like $f(R)$ (Carroll et al., 2004) or Galileon cosmology (Chow & Khoury, 2009). However, $\hat{\alpha}_{\mathrm{B/M}}<0$ can lead to gradient instabilities produced by an imaginary sound speed that causes exponentially growing perturbations (Zumalacárregui et al., 2017). Even on the theoretically stable $\hat{\alpha}_{\mathrm{M}}=-\hat{\alpha}_{\mathrm{B}}$ line, numerical fluctuations can lead to the aforementioned runaway effect. We demonstrate this effect by showing the $\hat{\alpha}_{\mathrm{M}}-\hat{\alpha}_{\mathrm{B}}$ parameter-space with successfully initialised hi_class instances in Fig. 4. Note that we use the prior ranges from Tab. 3 except for the braiding and Planck mass run rate, which we extend to $\hat{\alpha}_{\mathrm{B/M}}\in[-3,3]$. Thus, while certain sets of negative parameters are stable (Spurio Mancini et al., 2019), uniform and independent priors for $\hat{\alpha}_{\mathrm{B}}$ and $\hat{\alpha}_{\mathrm{M}}$ are only possible for positive values thereof, hence the choice in this work.
@@
To circumvent the issues of those projection affects, we additionally run our analysis for a “strong fiducial modified gravity” ($\hat{\alpha}_{\mathrm{B/M}}^{\mathrm{fid}}=1$) as opposed to the previously used “weak fiducial modified gravity” ($\hat{\alpha}_{\mathrm{B/M}}^{\mathrm{fid}}=0.05$). We know from the Planck CMB survey (Planck Collaboration et al., 2016) that we live at least in a close-to-$\Lambda$CDM Universe. Still, by increasing the fiducial braiding and Planck mass run rate parameters, we move them away from the prior edges and, hence, mimic looser boundaries. We do not quote the marginalised posterior uncertainties explicitly, but show the improvement when using FRBs in brackets in the last column of Tab. 2. However, we show the contours in Fig. 6. First, one should notice that much of the peculiar behaviour of the WL contours vanish in strong modified gravity; the WL $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$ posterior is no longer bimodal and the measurements of $k_{\mathrm{s}}$ are not biased towards lower values anymore. Consequently, the 68% confidence intervals change slightly as well. Even though FRBs still can not constrain modified gravity alone, the improvement of $\hat{\alpha}_{\mathrm{B}}$ from WL auto-correlation to WL-FRB cross-correlation slightly increases to $42.5\%$ and the one of $\hat{\alpha}_{\mathrm{M}}$ to $31.1\%$. Many of the other parameter improvements are comparable to the weak modified gravity case, except for $n_{\mathrm{s}}$ and $\sigma_{8}$ which respectively decrease and increase marginally in strong modified gravity. Therefore, we expect to gain constraining power on modified gravity when cross-correlating FRBs with WL, albeit less sizeable than for baryon-related parameters.
@@ 5 Application to FRBs with host
Figure 7: FRB with host identification used in this study. The error bars are a sum of $\sigma_{\mathrm{MW}}=50$, $\sigma_{\mathrm{host}}=50/(1+z)$ and the cosmological contribution from Eq. (36). The orange line corresponds to the best-fit model.
@@
where we use a $\delta$-distribution for the redshift distribution in Eq. (20) selecting redshift $z_{i}$ and $z_{j}$. Here the Legendre polynomials are denoted as $P_{\ell}(x)$.
@@
where $(\boldsymbol{\mathrm{DM}}_{\mathrm{host}})_{i}=\mathrm{DM}_{\mathrm{host}}/(1+z_{i})$ and $p_{\mathrm{LSS}}$ is a Gaussian with covariance with components given by Eq. (36) and the components of the mean by:
@@
In Fig. 7 we show the FRBs used for this analysis together with the best fit model. The error includes the covariance contribution from the LSS, from the host and from the Milky Way. Due to the low number of FRBs, we maximize the likelihood with respect to the $\Lambda$CDM parameters and then scan the profiles of $\hat{\alpha}_{\mathrm{M}}$ and $\hat{\alpha}_{\mathrm{B}}$ respectively. This removes the issue of projection effects and allows to get the most stringent constraints from the current data set. Fig. 8 shows this scan of the likelihood, showing that the posterior is prior driven (see Tab. 3). Particularly, there seems to be no information regarding $\hat{\alpha}_{\mathrm{M}}$ with the current data. While there is a slight constraint on $\hat{\alpha}_{\mathrm{B}}$, this is not significantly different from the prior. Generally this behaviour is expected as the parameter dependence is only present in the perturbations and thus in the covariance, degrading the information content compared to the model. However, it goes to show that host-identified FRBs are in principle a versatile tool to test gravity, in particular with many more sources expected in the next years (e.g. with DSA-2000, Hallinan et al., 2019)
Figure 8: Posterior distribution of $\hat{\alpha}_{\mathrm{B}}$ (blue) and $\hat{\alpha}_{\mathrm{M}}$ (orange) assuming a Gaussian (lognormal) host model in solid (dashed).
Parameter description fiducial value prior range source
$\log_{10}(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}})$ Horndeski theory screening scale -1 [$-2$, $0.3$] -
$h$ dimensionless Hubble parameter 0.674 [$0.38$, $1.00$] Planck Collaboration et al. (2020)
$\Omega_{\mathrm{b}}$ baryon density parameter $0.04931$ [$0.015$, $0.1$] Planck Collaboration et al. (2020)
$\Omega_{\mathrm{cdm}}$ cold dark matter density parameter $0.2642$ [$0.18$, $0.34$] Planck Collaboration et al. (2020)
$n_{\mathrm{s}}$ primordial scalar index $0.965$ [$0.7$, $1.25$] Planck Collaboration et al. (2020)
$\sum m_{\nu}/\mathrm{eV}$ total mass of all neutrino species $0.06$ [$0.003$, $1.5$] Planck Collaboration et al. (2020)
$\sigma_{8}$ linear power spectrum RMS fluctuations $0.811$ [$0.7$, $0.92$] Planck Collaboration et al. (2020)
fixed values $N_{\nu}$ number of neutrino species 3 Planck Collaboration et al. (2020)
$M_{\mathrm{Pl}}$ initial Planck mass $M_{\mathrm{Pl}}^{\mathrm{obs}}$
$\hat{\alpha}_{\mathrm{K}}$ kineticity function proportionality $1$
$N_{\mathrm{eff}}$ effective relativistic degrees of freedom 3.046 Planck Collaboration et al. (2020)
$T_{0}$ Black body CMB temperature today $2.7255\,\mathrm{K}$ Fixsen (2009)
$z_{\mathrm{reio}}$ reionisation redshift 7.7 Planck Collaboration et al. (2020)
$Y_{\mathrm{He}}$ primordial helium fraction 0.246 Planck Collaboration et al. (2020)
$w(a)$ dark energy equation of state -1 Planck Collaboration et al. (2020)
$\Omega_{\mathrm{K}}$ curvature density parameter 0 Planck Collaboration et al. (2020)
@@ Acknowledgements
During parts of this work, RR was supported by a European Research Council Consolidator Grant (No. 770935). DN acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101053992). SH was supported by the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311. RR is supported by the European Research Council (Grant No. 770935).
Appendix A Accuracy and details of the cosmopower emulator
Figure A.1: Histogram of the maximum and mean deviation of the emulated a) $P_{\mathrm{mm}}(k)$, b) $b_{\mathrm{e}}(k)$, c) $\eta(k)$, d) $\mu(k)$, e) $\chi(z)$ to the simulated test data.

### v2

@@ 3.1 MCMC likelihood
$\Omega_{\mathrm{b}}$ $0.044^{+0.006}_{-0.010}$ $0.049^{+0.011}_{-0.010}$ $0.0482^{+0.0029}_{-0.0025}$ 3.8 (3.8)
$\Omega_{\mathrm{cdm}}$ $0.240\pm 0.023$ $0.259\pm 0.012$ $0.2637\pm 0.0031$ 3.9 (3.3)
@@
3.2 Mock catalogue specification
3.2.1 FRB catalogue
We model the FRB catalogue properties according to a survey that is feasible within the next decade based on the predictions from Petroff et al. (2019). The actual number of host identified FRBs is rapidly growing. They are collected in the comprehensive Blinkverse-database [^9] https://blinkverse.zero2x.org/(Xu et al., 2023). As of the May 2025, there are 92 host identified FRBs. For this analysis, we choose $10^{4}$ FRB samples with associated redshift, whereas a substantial fraction is located above redshift 2. Assuming the FRBs trace the galaxy distribution (Reischke et al., 2021), the source distribution can be written as (assuming a flux limited sample)
$n(z)\propto z^{2}\exp(-\alpha z)\;,$ (32)
where $\alpha$ denotes the survey depth. Henceforth, we use $\alpha=2.5$ which fullfils the demands above. Note that the current sample of the CHIME survey roughly follows this shape with $\alpha=4$. However, several surveys are expected to be begin observations in the next years, amongst others the Canadian Hydrogen Observatory and Radio-transient Detector (CHORD, Vanderlinde et al. 2019) and the Deep Synoptic Array 2000 (DSA-2000, Hallinan et al. 2019). The latter one is expected to commence survey in March 2026 with an annual localised FRB detection rate of  104 and within 3.5 arcsec resolution, yielding a substantially larger detection rate and increased survey depth in comparison to CHIME, see Connor & Ravi (2023) for a more comprehensive review. We thus expect $\alpha=2.5$ to be more representative of the actual redshift distribution feasible within the next years. We split the redshift distribution into four equipopulated redshift bins (see also Tanidis & Camera 2019), thus recovering depths information and increasing the signal-to-noise ratio (SNR). We estimate the SNR from the Fisher information matrix (Tegmark et al., 1997) by treating the signal amplitude $A$ defined by $\hat{\boldsymbol{C}}_{\ell}=A\hat{\boldsymbol{S}}_{\ell}+\hat{\boldsymbol{N}}_{\ell}$ as a cosmological parameter, in which case the cumulative SNR can be calculated as
@@ 3.3 Three dimensional power spectrum generation
We calculate the linear matter power spectrum with the Boltzmann solver Horndeski in class (hi_class [^10] http://miguelzuma.github.io/hi_class_public/, Zumalacárregui et al., 2017, Bellini et al., 2020) suite which is an extension to the Cosmic Linear Anisotropy Solving System code (class [^11] https://github.com/lesgourg/class_public, Blas et al., 2011). Subsequently, utilising pyccl [^12] https://github.com/LSSTDESC/CCL (Chisari et al., 2019b), we pass the linear matter power spectrum into HMcode [^13] https://github.com/alexander-mead/HMcode (Mead et al., 2015) to calculate the non-linear matter and electron power spectrum $P_{\mathrm{mm}}$, $P_{\mathrm{\mathrm{ee}}}$ and, consequently, the electron bias via
@@
in Planck cosmology. The majority of signal resides around $\ell\sim\mathcal{O}(200)$ (see also Fig. 3) for FRBs. While cosmic shear obtains a significant signal from larger $\ell$, contributions for $k>50\;\mathrm{Mpc}$ will generally be small. For more details on the emulator choices we refer to appendix B.
@@ 4 Results
We show the 68% and 95% confidence regions for the three cases, i.e. WL and FRBs alone as well as WL $\times$ FRBs in Fig. 5. Overall, FRBs have considerably less constraining power than WL alone due to the overall lesser SNR, as WL from a stage-4 survey have a cumulative SNR which is roughly 10 times higher than the one of the FRB mock survey assumed here. Consequently, also the constraints on modified gravity are generally weaker. However, FRBs impose much smaller uncertainties on the baryon density $\Omega_{\mathrm{b}}h^{2}$ and the baryonic feedback strength $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$. The latter has been extensively discussed in our previous analysis (Reischke et al., 2023), hence we will refrain from detailing this effect here. Nevertheless, it should be noted that we expect a majority of the cross-correlation improvement to originate from degeneracy breaking due to baryonic feedback calibration with FRBs. Indeed, when cross-correlating shear and DM, the confidence regions of almost all parameters shrink significantly. The marginalised posterior mean and the 68% confidence intervals are quantified in Tab. 3. After $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$, the parameters describing the energy contents of the Universe ($\sum m_{\nu}$, $\Omega_{\mathrm{b}}$, $\Omega_{\mathrm{b}}$) experience the greatest increase in precision. Looking at the $\Omega_{\mathrm{b}}$-$h$ degeneracy apparent in Fig. 5, a separate calibration of $h$ further tightens the obtained constraints. In appendix A, we show the posterior for combining Euclid-CS×$\times$×FRB with the Planck2018 (Planck Collaboration et al., 2020a) temperature anisotropy auto-correlation. This breaks the degeneracy $\Omega_{\mathrm{b}}$-$h$, leading to much tighter constraints. Walters et al. (2018) similarly forecasted competitive constraining power when combining FRBs with distance ladder measurement on the background level. Since FRBs can be considered a late-type cosmological probe, this is a sensible addition for future analyses. A smaller but still considerable amount of gain is seen for the primordial power spectrum scale dependency $n_{\mathrm{s}}$, possibly caused by the comparatively large signal of FRBs for small $k$, at which $n_{\mathrm{s}}$ still determines the slope.
@@
Further peculiar behaviour can be seen for the $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$-Euclid-CS posterior in Fig. 5; displaying a bimodal distribution. While a Gaussian probability distribution cannot be expected in an MCMC, specifically the modified gravity parameter contour shapes show large deviations thereof caused by limiting prior ranges. We excluded significant influence from the MCMC-algorithm as the underlying cause (see App. C). Considering the strong deviation of the predicted $k_{\mathrm{s}}$ to the fiducial one, we conclude that the contours from Fig. 5 to be strongly influenced by the informative prior we place on $\hat{\alpha}_{\mathrm{B/M}}$ and $k_{\mathrm{s}}$.
@@
The motivation behind this choice is of both physical and numerical nature. In principle, $\hat{\alpha}_{\mathrm{B/M}}$ can be negative (Bellini & Sawicki, 2014) and even are expected to in many of the commonly investigated scalar-field theories like $f(R)$ (Carroll et al., 2004) or Galileon cosmology (Chow & Khoury, 2009). However, $\hat{\alpha}_{\mathrm{B/M}}<0$ can lead to gradient instabilities produced by an imaginary sound speed that causes exponentially growing perturbations (Zumalacárregui et al., 2017). Even on the theoretically stable $\hat{\alpha}_{\mathrm{M}}=-\hat{\alpha}_{\mathrm{B}}$ line, numerical fluctuations can lead to the aforementioned runaway effect. We demonstrate this effect by showing the $\hat{\alpha}_{\mathrm{M}}-\hat{\alpha}_{\mathrm{B}}$ parameter-space with successfully initialised hi_class instances in Fig. 4. Note that we simulate within the prior ranges from Tab. 2 except for the braiding and Planck mass run rate, which we extend to $\hat{\alpha}_{\mathrm{B/M}}\in[-3,3]$. Thus, while certain sets of negative parameters are stable (Spurio Mancini et al., 2019), uniform and independent priors for $\hat{\alpha}_{\mathrm{B}}$ and $\hat{\alpha}_{\mathrm{M}}$ are only possible for positive values thereof, hence the choice in this work. Additionally, we exclude $\hat{\alpha}_{\mathrm{B}}>2.5$ due to a decreased density of successful hi_class initialisations there as well.
Figure 7: FRB with host identification used in this study. The error bars are a sum of $\sigma_{\mathrm{MW}}=50$, $\sigma_{\mathrm{host}}=50/(1+z)$ and the cosmological contribution from Eq. 36. The orange line corresponds to the best-fit model.
Figure 8: Posterior distribution of $\hat{\alpha}_{\mathrm{B}}$ (blue) and $\hat{\alpha}_{\mathrm{M}}$ (orange) assuming a Gaussian (lognormal) host model in solid (dashed). The cutoff of $\hat{\alpha}_{\mathrm{B}}$ originates from prior limitations due to hi_class instabilities (see also Fig. 4).
@@
To circumvent the issues of those projection affects, we additionally run our analysis for a “strong fiducial modified gravity” ($\hat{\alpha}_{\mathrm{B/M}}^{\mathrm{fid}}=1$) as opposed to the previously used “weak fiducial modified gravity” ($\hat{\alpha}_{\mathrm{B/M}}^{\mathrm{fid}}=0.05$). We know from the Planck CMB survey (Planck Collaboration et al., 2016) that we live at least in a close-to-$\Lambda$CDM Universe. Still, by increasing the fiducial braiding and Planck mass run rate parameters, we move them away from the prior edges and, hence, mimic looser boundaries. We do not quote the marginalised posterior uncertainties explicitly, but show the improvement when using FRBs in brackets in the last column of Tab. 3. However, we show the contours in Fig. 6. First, one should notice that much of the peculiar behaviour of the WL contours vanish in strong modified gravity; the WL $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$ posterior is no longer bimodal and the measurements of $k_{\mathrm{s}}$ are not biased towards lower values anymore. The screening scale recovered now is on par with the fiducial input screening scale, chosen such as to suppress modified gravity in the non-linear regime (see also Spurio Mancini et al. 2018b).
Consequently, the 68% confidence intervals change slightly as well. Even though FRBs still can not constrain modified gravity alone, the improvement of $\hat{\alpha}_{\mathrm{B}}$ from WL auto-correlation to WL-FRB cross-correlation slightly increases to $42.5\%$ and the one of $\hat{\alpha}_{\mathrm{M}}$ to $31.1\%$. Many of the other parameter improvements are comparable to the weak modified gravity case, except for $n_{\mathrm{s}}$ and $\sigma_{8}$ which respectively decrease and increase marginally in strong modified gravity. Therefore, we expect to gain constraining power on modified gravity when cross-correlating FRBs with WL, albeit less sizeable than for baryon-related parameters. Adding CMB temperature anisotropy information has little effect on the modified gravity parameter contours, indicating only weak correlation between $\hat{\alpha}_{\mathrm{B/M}}$ and $h$, $\Omega_{\mathrm{b}}$, $n_{\mathrm{s}}$ (see appendix A).
@@ 5 Application to FRBs with host
where we use a $\delta$-distribution for the redshift distribution in Eq. 20 selecting redshift $z_{i}$ and $z_{j}$. Here the Legendre polynomials are denoted as $P_{\ell}(x)$.
@@
where $(\boldsymbol{\mathrm{DM}}_{\mathrm{host}})_{i}=\mathrm{DM}_{\mathrm{host}}/(1+z_{i})$ and $p_{\mathrm{LSS}}$ is a Gaussian with covariance with components given by Eq. 36 and the components of the mean by:
@@
In Fig. 7 we show the FRBs used for this analysis together with the best fit model, see Tab. 5 in the Appendix for a comprehensive list. The error includes the covariance contribution from the LSS, from the host and from the Milky Way. Due to the low number of FRBs, we maximize the likelihood with respect to the $\Lambda$CDM parameters and then scan the profiles of $\hat{\alpha}_{\mathrm{M}}$ and $\hat{\alpha}_{\mathrm{B}}$ respectively. This removes the issue of projection effects and allows to get the most stringent constraints from the current data set. Fig. 8 shows this likelihood scan, showing that the posterior is prior driven (see Tab. 2). Note that the sharp cut-off in $\hat{\alpha}_{\mathrm{B}}$ is caused by the introduced emulator boundaries. There seems to be no information regarding $\hat{\alpha}_{\mathrm{M}}$ with the current data. While there is a slight constraint on $\hat{\alpha}_{\mathrm{B}}$, this is not significantly different from the prior. Generally this behaviour is expected as the parameter dependence is only present in the perturbations and thus in the covariance, degrading the information content compared to the model. However, it goes to show that host-identified FRBs are in principle a versatile tool to test gravity, in particular with many more sources expected in the next years (e.g. with DSA-2000, Hallinan et al., 2019)
@@ Acknowledgements
During parts of this work, RR was supported by a European Research Council Consolidator Grant (No. 770935). DN acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101053992). SH was supported by the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311. RR is supported by the ERC (Grant No. 770935). H. Hildebrandt is supported by a DFG Heisenberg grant (Hi 1495/5-1), the DFG Collaborative Research Center SFB1491, an ERC Consolidator Grant (No. 770935), and the DLR project 50QE2305.
Appendix A Additional constraints from Planck CMB
We additionally investigate the influence of adding information from the Planck2018(Planck Collaboration et al., 2020a) temperature anisotropy power spectrum. Based on the chosen $\propto\Omega_{\Lambda}$ parametrisation used for modified gravity, we expect the main constraining power thereof to originate from degeneracy breaking between cosmological parameters. The posteriors are obtained by sampling the combined Planck2018-$\mathrm{TT}+$Euclid-CS×$\times$×FRB likelihood, whereas the Planck covariance is obtained from inverting the corresponding Fisher matrix and marginalising over nuisance parameters not included in the Euclid-CS×$\times$×FRB likelihood.
Figure A.1: 68% and 95 % posterior confidence regions (dark and light areas, respectively) for Euclid-CS×$\times$×FRB cross-correlation alone and in combination with Planck2018 temperature anisotropy auto-correlation. The fiducial cosmology taken from Tab. 2 is indicated by grey, dashed lines.
Furthermore, the Planck likelihood carries no information about the Horndeski screening scale and the baryonic feedback strength, hence we extend the Fisher matrix for $\log_{10}(\frac{T_{\mathrm{AGN}}}{\mathrm{K}})$ and $\log_{10}(\frac{k_{\mathrm{s}}}{\mathrm{Mpc}^{-1}})$ with small values ($10^{-5}$) along the corresponding diagonal, thus implying no intrinsic constraining power on these parameters. The posterior contours for the weak modified gravity case are shown in Fig. A.1. While there is a substantial increase in contraining power on $\Omega_{\mathrm{b}}$, the 68 % confidence region for $h$, and $n_{s}$, the $\hat{\alpha}_{\mathrm{B}}$ ($\hat{\alpha}_{\mathrm{M}}$) decreases only by a factor of 1.25 (1.17). The effect is even less for the strong modified gravity case, where the error margin decreases by a factor of 1.01 (1.04) for $\hat{\alpha}_{\mathrm{B}}$ ($\hat{\alpha}_{\mathrm{M}}$). Therefore, adding CMB temperature fluctuation information has only diminishing returns regarding modified gravity.
Appendix B Accuracy and details of the cosmopower emulator